The Roasted Bean charges $13.00 per pound for Fair Trade Organic Mexican coffee and $11.00 per pound for Fair Trade Organic Peruvian Coffee. How much of each type should be used to make a 28-lb blend that sells for $12.00 per pound?
Answer by stanbon(75887)

The Roasted Bean charges $13.00 per pound for Fair Trade Organic Mexican coffee and $11.00 per pound for Fair Trade Organic Peruvian Coffee. How much of each type should be used to make a 28-lb blend that sells for $12.00 per pound?
----
Quantity Eq: m + p = 28 pounds
Value Eq:: 13m+11p = 12*28 dollars
------------
Multiply thru the Quantity Eq. by 13:
13m + 13p = 13*28
13m + 11p = 12*28
----
Substact and solve for "p":
2p = 28
p = 14 lb (amt or Peruvian Coffee needed)
---
Sove for "m":
m + p = 28
m + 14 = 28
m = 14 lb (amt. of Mexican Coffee needed)
============================================
Cheers,
Stan H.
=========